OpenStudy (anonymous):

What does it mean to find two points that are (square root of 17) units apart?

OpenStudy (jdoe0001):

hmmm sounds pretty obvious :|

OpenStudy (jdoe0001):

what part is confusing you?

OpenStudy (anonymous):

Well were learning how to find the area of slanted squares and everybody was using 17 as the whole area. That didn't make sense to me

OpenStudy (jdoe0001):

well.... hmm what I can think of is....  \(\bf Area = length\cdot width=x\cdot x=x^2\\ \quad \\ \textit{if the area is 17, then what's "x"?}\\ \quad \\ Area = 17\qquad Area = x^2\qquad x^2=17\implies x=\sqrt{17}\)

OpenStudy (anonymous):

oh! That makes sense. Your right. Stupid question

OpenStudy (jdoe0001):

\(\bf Area = x^2\implies Area = \sqrt{17}\cdot \sqrt{17}\implies Area=\sqrt{17^2}\implies Area = 17\)